A piece of wire 7 / 8 meter long broke into two pieces. One-piece was 1 / 4 meter long. How long is the other piece?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Total length of wire = 7 / 8 metre
Length of one piece of wire = 1 / 4 metre
Length of other piece of wire = Length of the original wire and this one piece of wire
=7 / 8–1 / 4
= [(7× 1)–(1 × 2)] / 8
= (7–2) / 8
=5 / 8
∴Length of the other piece of wire = 5 / 8 metre

​Thanks